OpenStudy (anonymous):

If m angle 2= 7x + 7, m angle 3 = 5y, and m angle 4= 140, find the values of x and y. x = 19, y = 8 x = 140, y = 40 x = 8, y = 19 x = 40, y = 140

OpenStudy (anonymous):

Help mee

OpenStudy (mathstudent55):

Is there a figure that shows the relationship between the angles?

OpenStudy (anonymous):

yup

OpenStudy (mathstudent55):

Ok. Do you know the name of angles 1 and 3 in the figure below?

OpenStudy (anonymous):

vertical angles ?

OpenStudy (mathstudent55):

Correct. What do you know about the measures of vertical angles?

OpenStudy (anonymous):

im about to get bumped /:

OpenStudy (anonymous):

they are equal

OpenStudy (mathstudent55):

Correct. Two more questions about the angles. WHat are angles 1 and 2 called, and what do you now about their measures?

OpenStudy (anonymous):

they are supplementary, 180

OpenStudy (mathstudent55):

Wow, you know your stuff. Great. Now we are ready to solve this problem.

OpenStudy (anonymous):

auuuuright

OpenStudy (mathstudent55):

Angles 1 and 4 are vertical and have the same measure. m<2 = 7x + 7 m<4 = 140 That means 7x + 7 = 140 Now we solve that equation.

OpenStudy (mathstudent55):

Do you know how to solve that equation?

OpenStudy (mathstudent55):

You need to isolate x. First, subtract 7 from both sides. Then divide both sides by 7.

OpenStudy (anonymous):

19 ?

OpenStudy (mathstudent55):

Correct. x = 19 Now we need y.

OpenStudy (mathstudent55):

We know this about y: m<3 = 5y We need to find the measure of angle 3, then we can find y.

OpenStudy (mathstudent55):

Angle 3 is supplementary to angle 4. The measure of angle 4 is 140. What is the measure of angle 3?

OpenStudy (anonymous):

8 ?

OpenStudy (anonymous):

thanks! @mathstudent55

OpenStudy (mathstudent55):

Correct. m<3 is 40 3y = 40 y = 8 Great job!

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (anonymous):

If m<2 = 7x + 7, m<3 = 4y, and m<4 = 112, find the values of x and y. x = 15, y = 17 x = 112, y = 68 x = 17, y = 15 x = 68, y = 112 @mathstudent55

OpenStudy (anonymous):

i need help agen

OpenStudy (mathstudent55):

Same figure?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

Measures of angles 2 and 4 are the same since angles 2 and 4 are vertical. Set their measures equal and solve for x.

OpenStudy (mathstudent55):

What do you get for x?

OpenStudy (anonymous):

x=15 and y=17 ?! (: right ??

OpenStudy (mathstudent55):

Correct. Great job again!

OpenStudy (mathstudent55):

gtg, bye